# promptbox[1][] colframe=black, colback=white, sharp corners, boxrule=1pt, boxsep=8pt, left=0pt, right=0pt, top=0pt, bottom=0pt, #1

### v1

@@ 4.2.2. Personas
Persona ID Age Gender Education level Race/Ethnicity Self Rated Digital Literacy
@@
3 20 Male High school diploma White/Caucasian 5
@@
5 25 Male Bachelor Black/African American 5
6 28 Female Master White/Caucasian 5
7 23 Female Master Hispanic/Latino 4
@@
Table 1. The demographic information of real personas. The digital literacy was rated on a 5 Likert scale where 1 stands for “not at all proficient” and 5 represents “highly proficient”.
@@
{promptbox}
[title=Prompt to generate social media post content for GPT-generated personas] Provide ideas for this person to write posts (limit the word to 140 words).
Output the posts in the following JSON format in plain text: { “time”: ¡time in string format¿, “address”: ¡address where this person shares the life¿, “content”: ¡content¿, “latitude”: ¡fake latitude¿, “longitude”: ¡fake longitude¿, “timezone”: ¡time zone¿, “locale”: ¡locale¿ }
@@ 4.2.3. Websites
Website Topics Number of Advertisements Site Rank
www.weather.com Weather forecasts and news 8 37
@@
www.researchgate.net Academic articles 3 556
www.usnews.com National news, college rankings 3 1,165
www.fashionista.com Fashion and celebrity news 5 78,490
@@ 4.3. Data analysis methods
To analyze the quantitative data gathered from the Likert scale survey, we employed one-way ANOVA for each survey item to evaluate the significant difference in mean scores regarding the personas’ realness among three groups: our approach, the baseline GPT, and real personas. Whenever significant differences emerged, we conducted post-hoc tests using Tukey’s pairwise comparisons to gain deeper insights into these distinctions.
For the qualitative data analysis, we followed established open coding procedures (Brod et al., 2009). Two members of our research team independently initiated the coding process in MAXQDA. One researcher coded 20 percent of the sample and generated a set of initial codes. Subsequently, the second researcher coded the same portion to introduce new codes if necessary. Non-agreement cases were discussed to reconcile differences and establish a cohesive codebook. Utilizing this codebook, we conducted a thematic analysis to uncover and delineate the significant themes that emerged during the interviews and were pertinent to the established codes. The complete codebook is presented in the Appendix C. These themes were then consolidated and evolved into study findings that are detailed in Section 5.
@@ 5. Results and Findings
5.1. Users’ Perceived Realness of Privacy Personas
@@
We applied one-way ANOVA and posthoc tests to analyze the difference in perceived realness of users for three groups of personas. Fig. 6 shows the mean and standard error for each measure, including credibility, consistency, clarity, and empathy.
@@ 5.1.4. Empathy:
Our results suggest that personas created using our method improve users’ understanding of the personas’ motivations when compared to those generated by GPT. However, the behavior of real personas is influenced by intricate factors. This complexity may cause users to exhibit slightly reduced cognitive empathy for real personas as compared to those we generated. Interestingly, users showed no significant difference in emotional empathy across the three persona categories. The overall empathy scores (both cognitive and emotional) were moderately low. This could be because users only reviewed the profiles and did not immerse themselves in the personas’ identities to experience the impact of privacy attributes on system outcomes. However, we observed that users expressed noticeable excitement or surprise when using personas’ identities and encountering highly relevant advertisements (details in Section 5.4). This suggests that relying solely on browsing personas’ information has limited efficacy in eliciting empathy. The actual interactive experience of using personas through the Privacy Sandbox might be necessary to foster greater empathy towards personas.
5.2. Factors Influencing the Perceived Realness of Generated Personas
@@ 5.2.2. Deviation from personal experiences
    Specific Attributes: Participants questioned a persona’s realness if certain details did not align with their experiences. For instance, after seeing personas of a financial analyst and a cashier, six participants felt the given income was too low. A participant noted, “One thing that I would notice is usually financial analysts make a lot of money. So \$70,000, this annual income does not seem reasonable. This seems unusual to me, given that Michael was born in 1981, he is 40 years old, so he probably has a lot of experience in the field. He definitely should have been making more than \$70,000.” Another issue raised was about browsing history; participants felt the content was too basic for an experienced individual, with one commenting, “Given his age, I think he is an experienced custodian, does not need to search for the information about this job itself.
@@
This pertains to inconsistencies in specific details when compared to the participants’ own experiences. Discrepancies were noticed in weekly schedules, browsing histories, and social media posts.
@@
    Weekly Schedule: All participants felt the work hours for the personas were too lengthy, with observations such as, “It looks like they work at least 8 hours a day. Notice that I mean yeah 8 hours a day for seven days a week.
@@
    Browsing History: Four participants believed the browsing history was too centered on work, lacking diversity. One commented, “The website she is browsing is all about her job uh is all related to her job and her professional, but uh they are. There should be some other content about her life.” They also expected more continuity in browsing, mentioning, “I would continue to click the content in those websites. So again, those links, those four links should be the same…should be consistent or should be progressive.” Participants were also skeptical of identical timestamps on different records, with a remark, “She cannot be on the same page at the exact same time. This is to the second exact same time. This is incorrect.
@@
    Social Media Posts: Some participants felt the posts were overly positive and superficial. One observation was, “I feel strange about his posts is that he always appears so positive, like his life is so perfect and he’s very proud of his work.” They thought that the contents were superficial, and lacked depth in emotions and thoughts, as one participant mentioned, “There’s no real emotion…it just feels like she wants to prove something to you.
@@ 5.2.3. Consistency within the data
    Out-of-Context Content: Users identified certain generated browsing histories as not aligning with the persona’s description. For instance, “In the afternoon and evening, he’s not using the internet or his mobile phone. And in the schedule, it should have more history about liking YouTube videos, social media interaction.” Such discrepancies were attributed to sample leakage during the few-shot learning process. For example, the model was exposed to browsing samples related to bike garages. To enhance the realness and relevance of generated data, we must address such technical issues in future iterations.
5.3. System Outcomes Influenced by Privacy Data Replacement
@@
$ad\ overlap\ rate=\frac{number\ of\ duplicated\ ads\ between\ personas}{total\ number\ of\ ads\ in\ a\ website}$
@@
For every website considered, we began by gathering all the advertisements presented when accessing the site with each of the eight personas. Among these ads, we specifically noted those that appeared for multiple personas—effectively highlighting the number of overlapping or duplicate ads across personas.
Website duplicated ads numbers total ads numbers ad overlap rate
@@
Figure 7. Examples of different ads encountered by participants when browsing the selected websites using various persona identities. Note that participants using the first persona received ads about shows in Atlanta (associated with the location), investment tools (associated with the profession and browsing history), and shoes for men (associated with gender) while participants using the second persona received different ads associated with her attributes.
@@
Table 3 shows the result of the ad overlap rate for each selected website. The overlap rates for all websites are less than $50\%$. This implies that when users switch between eight personas, over half of the ads they encounter are unique to a single persona. Fig. 7 shows the variations in ads on weather.com for different personas. In subsequent sections, we will dive deeper into how users associate these advertisements with the underlying private information of the personas.
5.4. Perceived Links between Privacy Data and System Outcomes
@@
All participants, while browsing the websites as personas, encountered ads related to the persona’s private attributes. Participants perceived such connections in two ways.
@@
(1) Direct connection based on privacy attributes: Participants noticed an explicit correspondence between the ads and the persona’s privacy attributes. Participants often associated ads with personas based on their interests, hobbies, daily activities, occupation, educational background, marital status, and family situation. For example, one participant claimed that an Xfinity ad was relevant to the selected persona because “he does a lot of social media, gaming, YouTube. So he maybe wants to use Xfinity to watch something.” When an ad seemed especially pertinent to a persona, it evoked reactions of excitement or surprise. For instance, when a participant was browsing the website through a persona living in Los Angeles and saw an ad promoting environmental protection in the same area, she said, “[The ad is] very real because it’s located in Los Angeles.” Sometimes, even when users subjectively did not consider an advertisement to be highly related to the persona, they could still speculate about the reasons for encountering the ad. A participant said “Max may be tangentially related because he is interested in gaming. So the algorithm might assume that somebody who’s interested in gaming might be interested in media services as well.” These reasoning processes and emotional reactions showcase participants’ cognitive and affective empathy toward personas. Furthermore, it underscores their enhanced privacy knowledge through the examination of ad-persona relationships.
@@
(2) Indirect connection based on personal association and stereotypes: Interestingly, sometimes users could not directly pinpoint a specific privacy attribute related to an ad, but they still considered the ad relevant to the persona because they made associations based on known privacy information about the persona. Such judgments were sometimes made on their personal biases or societal stereotypes. A typical example is when a participant believed that a tire ad was related to an African-American persona. Although the persona’s profile did not mention any information related to tires, the participant expressed, “I think honestly the [ad] save time and money for tires is more for African Americans. Because of African American culture, they like to modify their tires.
@@ 6. Discussion
6.1. A trade-off in the impact of familiarity and persona realness on empathy
@@
Ideally, users should view personas as both familiar and authentic. Yet, the constraints of current large language models hinder achieving optimal realness (Hämäläinen et al., 2023). One method to compensate for the impact of reduced realness is to maintain a balance in familiarity. The goal is to prevent users from seeing personas as too alien or too familiar, which might reveal data inaccuracies. To find the equilibrium and explore the trade-off, we intend to conduct more in-depth user studies, allowing users to experience generated personas with varying levels of realness and familiarity. These studies will help us assess users’ empathy towards these personas and delineate the interrelation between familiarity and perceived realness in invoking empathy.
6.2. The influence of personal views on empathy and privacy literacy
User empathy towards personas is influenced not only by their familiarity with the domain but also by their personal views. As highlighted in Section 5.2.2, participants with different personal views perceived the privacy attributes of the same persona differently. Deviations between a persona’s privacy attributes and a user’s views can lower the persona’s perceived realness. Such views can arise from personal experiences, observations, or even stereotypes. The personal views of users may stem from their own experiences (e.g., a participant working in the finance sector feeling that the data for a “fanatical analyst” persona is not realistic enough), observations of others’ experiences (e.g., a participant with research assistant friends believing that research assistants should have longer working hours), or even personal stereotypes (e.g., a participant thinking that African Americans enjoy changing tires).
@@ 6.2. The influence of personal views on empathy and privacy literacy
6.3. Acquisition of privacy knowledge
@@
This demonstrates the potential of using our proposed approach to support users in experiential learning (Gentry, 1990) for privacy knowledge. Our privacy sandbox prototype provides users with personal involvement in learning privacy knowledge, as both the feelings and cognitive aspects of users are engaged. Furthermore, this approach serves as a way of scaffolding (Gonulal and Loewen, 2018), enabling users to experience the influence of privacy attributes on system outcomes that they were unable to manage on their own (Miller, 2002).
Through our approach, users can independently contextualize the privacy attributes of specific personas and can understand how the system uses privacy information for personalized content. To validate the effectiveness of this approach in enhancing users’ acquisition of privacy knowledge, future steps involve using rigorous tests to assess users’ privacy knowledge before and after using the approach. Furthermore, observing how users apply the privacy knowledge they gain from this approach in simulated or real-world settings can provide deeper insights into how the knowledge impacts users’ privacy behaviors.
@@ 6.4. Biases in LLMs and their impact on generated personas
Previous work has shown that Large Language Models (LLMs) like GPT risk amplifying existing stereotypes (Blodgett et al., 2020; Nadeem et al., 2020; Smith et al., 2022). However, these biases within the personas generated may not be detrimental in the specific context of our study.
@@ 6.6.1. Generation pipeline
While our current generation pipeline was shown to be generally capable of creating artificial personas that are sufficiently realistic to stimulate changes in system outcomes and invoke user empathy, in Section 5.2.3, users pointed out issues of inconsistencies in currently generated persona data. We found that these inconsistencies were partly due to the inherent randomness embedded in the output of large language models and also resulted from inductive biases (Lake et al., 2019) during the few-shot learning process. Such biases can cause generated data to resemble the example data, resulting in users perceiving the generated information as out of context. Our planned future steps to address this issue include fine-tuning the model to enhance data consistency and reduce the generation of irrelevant data. We also plan to store and represent the facts generated about artificial personas more formally in a knowledge graph and improve the coherence within the generated data using knowledge infusion techniques (Moiseev et al., 2022; Heinzerling and Inui, 2020).

### v2

@@ 4.2.2. Personas
ID Age Gender Education level Ethnicity Digital Literacy
@@
3 20 Male High school White 5
@@
5 25 Male Bachelor Black 5
6 28 Female Master White 5
7 23 Female Master Hispanic 4
@@
Table 1. The demographic information of real personas. The digital literacy was self rated on a 5 Likert scale where 1 stands for “not at all proficient” and 5 represents “highly proficient”.
@@
Prompt to generate social media post content for GPT-generated personas Provide ideas for this person to write posts (limit the word to 140 words). Output the posts in the following JSON format in plain text: { “time”: ¡time in string format¿, “address”: ¡address where this person shares the life¿, “content”: ¡content¿, “latitude”: ¡fake latitude¿, “longitude”: ¡fake longitude¿, “timezone”: ¡time zone¿, “locale”: ¡locale¿ }
@@ 4.2.3. Websites
Website Topics # of Ads Site Rank
www.weather.com Weather forecasts 8 37
@@
www.researchgate.net Academic paper 3 556
www.usnews.com National news 3 1,165
www.fashionista.com Fashion 5 78,490
@@ 4.3. Data analysis methods
To analyze the quantitative data collected from the Likert scale survey, we employed one-way ANOVA for each survey item to evaluate the significant difference in mean scores regarding the personas’ realness among three groups: our approach, the baseline GPT, and real personas. Whenever significant differences emerged, we conducted post-hoc tests using Tukey’s pairwise comparisons to gain deeper insights into these distinctions.
For qualitative data analysis, we followed established open coding procedures (Brod et al., 2009). Two members of our research team independently initiated the coding process in MAXQDA. A researcher coded 20% of the sample and generated a set of initial codes. Subsequently, the second researcher coded the same portion to introduce new codes if necessary. Non-agreement cases were discussed to reconcile differences and establish a cohesive codebook. Utilizing this codebook, we conducted a thematic analysis to uncover and delineate the significant themes that emerged during the interviews and were relevant to the established codes. The complete codebook is presented in Appendix C. These themes were then consolidated and evolved into study findings that are detailed in Section 5.
@@ 5. Results and Findings
5.1. Users’ perceived realness of privacy personas (RQ1)
@@
We applied one-way ANOVA and post-hoc tests to analyze the difference in perceived realness of users for three groups of personas. Fig. 6 shows the mean and standard error for each measure, including credibility, consistency, clarity, and empathy.
@@ 5.1.4. Empathy:
Our results suggest that personas created using our method improve users’ understanding of the personas’ motivations when compared to those generated by GPT. However, the behavior of real personas is influenced by intricate factors. This complexity may cause users to exhibit slightly reduced cognitive empathy for real personas compared to those we generated. Interestingly, users showed no significant difference in emotional empathy across the three persona categories. The overall empathy scores (both cognitive and emotional) were moderately low. This could be because users only reviewed the profiles and did not immerse themselves in the personas’ identities to experience the impact of privacy attributes on system outcomes. However, we observed that users expressed noticeable excitement or surprise when using personas’ identities and encountering highly relevant advertisements (details in Section 5.4). This suggests that relying solely on browsing personas’ information has limited efficacy in eliciting empathy. The actual interactive experience of using personas through the Privacy Sandbox might be necessary to foster greater empathy towards personas.
5.2. Factors influencing the perceived realness of generated personas (RQ2)
@@ 5.2.2. Deviation from personal experiences
    Specific Attributes: Participants questioned the realness of a persona if certain details did not align with their experiences. For instance, after seeing personas of a financial analyst and a cashier, six participants felt the given income was too low. A participant noted, “One thing that I would notice is usually financial analysts make a lot of money. So \$70,000, this annual income does not seem reasonable. This seems unusual to me, given that Michael was born in 1981, he is 40 years old, so he probably has a lot of experience in the field. He definitely should have been making more than \$70,000.” Another issue raised was about browsing history; participants felt the content was too basic for an experienced individual, with one commenting, “Given his age, I think he is an experienced custodian, does not need to search for the information about this job itself.
@@
This pertains to inconsistencies in specific details when compared to the participants’ own experiences. Discrepancies were observed in weekly schedules, browsing histories, and social media posts.
@@
    Weekly Schedule: All participants felt that the work hours for the personas were too lengthy, with observations such as, “It looks like they work at least 8 hours a day. Notice that I mean yeah 8 hours a day for seven days a week.
@@
    Browsing History: Four participants believed that browsing history was too centered on work, lacking diversity. One commented, “The website she is browsing is all about her job uh is all related to her job and her professional, but uh they are. There should be some other content about her life.” They also expected more continuity in browsing, mentioning, “I would continue to click the content in those websites. So again, those links, those four links should be the same…should be consistent or should be progressive.” Participants were also skeptical of identical timestamps on different records, with a remark, “She cannot be on the same page at the exact same time. This is to the second exact same time. This is incorrect.
@@
    Social Media Posts: Some participants felt that the posts were overly positive and superficial. One observation was, “I feel strange about his posts is that he always appears so positive, like his life is so perfect and he’s very proud of his work.” They thought that the contents were superficial, and lacked depth in emotions and thoughts, as one participant mentioned, “There’s no real emotion…it just feels like she wants to prove something to you.
@@ 5.2.3. Consistency within the data
    Out-of-Context Content: Users identified certain generated browsing histories as not aligning with the persona’s description. For instance, “ there are many histories about bike riding, but in his profile, there is no information about riding a bike.” Such discrepancies were attributed to sample leakage during the few-shot learning process. For example, the model was exposed to browsing samples related to bike garages. To enhance the realness and relevance of generated data, we must address such technical issues in future iterations.
5.3. System outcomes influenced by privacy data replacement (RQ3)
@@
ad overlap rate=number of duplicated ads between personastotal number of ads in a $\text{ad overlap rate}=\frac{\text{number of duplicated ads between personas}}{\text{total number of ads in a website}}$
@@
For every website considered, we began by gathering all the advertisements presented when accessing the site with each of the eight personas. Among these ads, we specifically noted those that appeared for multiple personas, effectively highlighting the number of ads that overlap or duplicate across personas.
Website duplicated ads total ads ad overlap rate
@@
Figure 7. Examples of different ads encountered by participants when browsing the selected websites using various persona identities. Note that participants who used the first persona received ads about shows in Atlanta (associated with the location), investment tools (associated with the profession and browsing history), and shoes for men (associated with gender), while participants who used the second persona received different ads associated with her attributes.
@@
Table 3 shows the result of the ad overlap rate for each selected website. The overlap rates for all websites are less than $50\%$. This implies that when users switch between eight personas, more than half of the ads they encounter are unique to a single persona. Fig. 7 shows the variations in the ads on weather.com for different personas. In subsequent sections, we will dive deeper into how users associate these advertisements with the underlying private information of the personas.
5.4. Perceived links between privacy data and system outcomes (RQ4)
@@
All participants, while browsing the websites as personas, encountered ads related to the persona’s privacy attributes. When inquired about the ads, participants perceived the connections between the ads and the persona in two ways.
@@
(1) Direct connection based on privacy attributes: Participants noticed an explicit correspondence between the ads and the persona’s privacy attributes. Participants often associated ads with personas based on their interests, hobbies, daily activities, occupation, educational background, marital status, and family situation. For example, one participant claimed that an Xfinity ad was relevant to the selected persona because “he does a lot of social media, gaming, YouTube. So he maybe wants to use Xfinity to watch something.” When an ad seemed especially pertinent to a persona, it evoked reactions of excitement or surprise. For instance, when a participant was browsing the website through a persona living in Los Angeles and saw an ad promoting environmental protection in the same area, she said, “[The ad is] very real because it’s located in Los Angeles.” Sometimes, even when users subjectively did not consider an advertisement to be highly related to the persona, they could still speculate about the reasons for encountering the ad. A participant said “Max may be tangentially related because he is interested in gaming. So the algorithm might assume that somebody who’s interested in gaming might be interested in media services as well. These reasoning processes and emotional reactions demonstrate participants’ cognitive and affective empathy toward personas. Furthermore, it underscores their enhanced privacy knowledge through the examination of ad-persona relationships. By summarizing and generalizing the patterns evolved from the correlations, users may leverage such knowledge to reflect on and improve their privacy settings and configurations.
@@
(2) Indirect connection based on personal association and stereotypes: Interestingly, sometimes users were unable to directly pinpoint a specific privacy attribute related to an ad, but still considered the ad relevant to the persona because they made associations based on known privacy information about the persona. Such judgments were sometimes made on their personal biases or societal stereotypes. A typical example is when a participant believed that a tire ad was related to an African-American persona. Although the persona’s profile did not mention any information related to tires, the participant expressed, “I think honestly the [ad] save time and money for tires is more for African Americans. Because of African American culture, they like to modify their tires.
@@ 6. Discussion
6.1. A trade-off in the impact of familiarity and persona realness on empathy (RQ2)
@@
Ideally, users should view personas as both familiar and authentic. Yet, the constraints of current large language models hinder achieving optimal realness (Hämäläinen et al., 2023). One method of compensating for the impact of reduced realness is to maintain a balance in familiarity. The goal is to prevent users from seeing personas as too alien or too familiar, which might reveal data inaccuracies. To find the equilibrium and explore the trade-off, we intend to conduct more in-depth user studies, allowing users to experience generated personas with varying levels of realness and familiarity. These studies will help us assess users’ empathy towards these personas and delineate the interrelation between familiarity and perceived realness in invoking empathy.
6.2. The influence of personal views on empathy and privacy literacy (RQ2 & RQ4)
User empathy towards personas is influenced not only by their familiarity with the domain, but also by their personal views. As highlighted in Section 5.2.2, participants with different personal views perceived the privacy attributes (e.g., income, weekly schedule, and browsing history) of the same persona differently. Deviations between a persona’s privacy attributes and a user’s views can lower the persona’s perceived realness. Such views can arise from personal experiences, observations, or even stereotypes. The personal views of users may stem from their own experiences (e.g., a participant working in the finance sector feeling that the data for a “financial analyst” persona is not realistic enough), observations of others’ experiences (e.g., a participant with research assistant friends believing that research assistants should have longer working hours), or even personal stereotypes (e.g., a participant thinking that African Americans enjoy changing tires).
@@ 6.2. The influence of personal views on empathy and privacy literacy (RQ2 & RQ4)
6.3. Acquisition of privacy knowledge (RQ3 & RQ4)
@@
This demonstrates the potential of using our proposed approach to support users in experiential learning (Gentry, 1990) for privacy knowledge. Our privacy sandbox prototype provides users with personal involvement in learning privacy knowledge, as both the feelings and the cognitive aspects of users are engaged. Furthermore, this approach serves as a way of scaffolding (Gonulal and Loewen, 2018), enabling users to experience the influence of privacy attributes on system outcomes that they were unable to manage on their own (Miller, 2002).
Through our approach, users can independently contextualize the privacy attributes of specific personas and understand how the system uses privacy information for personalized content. To validate the effectiveness of this approach in enhancing users’ acquisition of privacy knowledge, future steps involve using rigorous tests to assess the privacy knowledge of users before and after using the approach. Furthermore, observing how users apply the privacy knowledge they gain from this approach in simulated or real-world settings can provide deeper insights into how the knowledge impacts users’ privacy behaviors.
@@ 6.4. Biases in LLMs and their impact on generated personas
Previous work has shown that Large Language Models (LLMs) like GPT risk amplifying existing stereotypes (Blodgett et al., 2020; Nadeem et al., 2020; Smith et al., 2022; Shen et al., 2023). However, these biases within the personas generated may not be detrimental in the specific context of our study.
@@ 6.6.1. Generation pipeline
While our current generation pipeline was shown to be generally capable of creating artificial personas that are sufficiently realistic to stimulate changes in system outcomes and invoke user empathy, in Section 5.2.3, users pointed out the issues of inconsistencies in the current generated persona data. We found that these inconsistencies were partly due to the inherent randomness embedded in the output of large language models and also resulted from inductive biases (Lake et al., 2019) during the few-shot learning process. Such biases can cause generated data to resemble the example data, resulting in users perceiving the generated information as out of context. Our planned future steps to address this issue include fine-tuning the model to enhance data consistency and reduce the generation of irrelevant data. We also plan to store and represent the facts generated about artificial personas more formally in a knowledge graph and improve the coherence within the generated data using knowledge infusion techniques (Moiseev et al., 2022; Heinzerling and Inui, 2020).